OpenStudy (falloutboy):

A circle has a radius of 8 inches. Find the area of a sector of the circle if the sector has an arc that measures 45°. a)2 sq. in. b)8 sq. in. c)16 sq. in.

OpenStudy (owlcoffee):

You can find the area of any sector in a circle, by having the angle and the radius and use the followinf formula: \[A _{s \circ }=(\frac{ \alpha }{ 360 })(\pi r^2)\] Where alpha is the angle and "r" is the radius.

OpenStudy (falloutboy):

...???

OpenStudy (owlcoffee):

just replace the values in the formula: \[A=(\frac{ 45 }{ 360 })(\pi (8)^2 )\]

OpenStudy (falloutboy):

oh ok....sorry ive never done this befor...and thanks